OpenStudy (anonymous):

solve linear system using and algebraic method 3x+ 3y = 3 y=-x

OpenStudy (anonymous):

\[3x+3(-x)=3\]\[3(-y)+3y=3\] Then simplify.

OpenStudy (radar):

3x-3x=3 0=3 ????? No solutions

OpenStudy (anonymous):

wait so whats the right answer?

OpenStudy (radar):

The right answer is there is no answer. I guess you could say y can not be equal to -x.

OpenStudy (anonymous):

if \[x=-y\] then \[3x+3y=0\]

OpenStudy (anonymous):

for example if \[x=4\] then \[y=-4\] and \[3\times 4+3\times -4=12-12=0\] so you cannot get 3 out of it

OpenStudy (anonymous):

It's a system with no solution, since 3 cannot equal zero.

OpenStudy (anonymous):

okay thank you